# Automated Knowledge Concept Annotation and Question Representation Learning for Knowledge Tracing

### v1

@@ Appendix F Performance on Multi-step ahead Prediction Task
Fig. 10 demonstrates the extended prediction results for KT models for multi-step ahead prediction task with both scenarios: (a) accumulative and (b)  non-accumulative prediction. Of note, we excluded IEKT from this task due to its slow inference. As described in the implementation details (Appendix D), our KCQRL does not affect the models’ original runtimes, so the slow inference is solely attributed to IEKT’s original implementation.
@@
Figure 10: Improvement of our KCQRL across models and datasets in multi-step ahead prediction scenario. Green [red] area covers the improvement [decline] from our framework.
Appendix G Quality of Automated KC Annotation
In this section, we provide the details how we evaluate the quality of our KCQRL’s KC annotations as part of our ablation study in Sec. 5.2. For this, we compared our annotations against two baselines: (1) KC annotations from the original datasets and (2) KC annotations of our KCQRL without leveraging the solution steps.
@@ Appendix G Quality of Automated KC Annotation
For a structured comparison, we introduce the following five criteria. (1) Correctness measure which annotation has correct KC(s), relative to the other annotations. (2) Coverage is an evaluation criterion for questions with multiple KCs. It measures how well the KCs from an annotation cover the set of KCs that the question is associated with. (3) Specificity compares which annotation has a more modular set of KCs rather than long and complicated ones. Of note, modular and simpler KCs enable identifying the common skills required for different questions. (4) We additionally defined ability of integration to evaluate how well the described KCs are widely applicable to other Math problems, beyond the question being solved. (5) Finally, overall is about choosing the best KC annotation by considering all criteria.
As explained in Sec. 5.2, we made a pair-wise comparison between KC annotations: (i) original vs. KCQRL w/o sol. steps; (ii) original vs. KCQRL,; and (iii) KCQRL w/o sol. steps vs. KCQRL. To eliminate any potential bias from the order of KC annotations in the prompt, we randomly assigned KC annotations to groups A and B. Our prompt is given in Appendix G.2.
G.1 Example KC Annotations
Here, we build upon our example question from Fig. 2 and provide the KC annotations of two baselines and our framework.
{mdframed}
@@ G.1 Example KC Annotations
[…]
$\Rightarrow$ Takeaway: The original KC annotation is just one phrase with complex combinations of multiple KCs. It is also missing out some KCs such as addition and simplifications of expressions. On the other hand, KC annotations of KCQRL w/o solution steps are missing the important techniques asked by the problem, such as extracting the common factor. The reason is that this technique is hidden in the solution steps, which thus need to be provided for a better KC annotation. Overall, our complete framework correct set of KCs with better coverage and in a modular way in comparison to two baseline annotations.
G.2 Prompt for Quality Comparison of KC Annotations
@@
Appendix H Prompts for KC Annotation via LLMs
@@
H.1 Solution step generation
@@ H.2 KC Annotation
H.3 Solution Step-KC Mapping
As the final part, our framework maps each solution step to its associated KCs for a given question. This step is particularly needed for Module 2 of our framework: representation learning of questions (Sec. 3.2). The prompt of this step is given below.

### v2

@@ Appendix F Performance on Multi-step ahead Prediction Task
Fig. 10 demonstrates the extended prediction results for KT models for multi-step ahead prediction task with both scenarios: (a) accumulative and (b)  non-accumulative prediction. Of note, we excluded IEKT from this task due to its slow inference. As detailed in the implementation section (Sec.D), our KCQRL does not affect the models’ original runtimes, so the slow inference is solely attributed to IEKT’s original implementation.
@@
Figure 10: Improvement of our KCQRL across models and datasets in multi-step ahead prediction scenario. Green/red area covers the improvement/decline from our framework.
Appendix G Visualization of Question Embeddings (Extended)
Fig. 11 plots the question embeddings for ablations and compares them against our complete KCQRL framework. Specifically, we have five ablations without any CL training from the same LM encoder: (a) question content, (b) question and its KCs, (c) question and its solution steps, (d) question and its solution steps and KCs, and (e) only the KCs of a question. Further, to demonstrate the effectiveness of our CL loss, we implement the following ablations: (f) KCQRL w/o false negative elimination, whose CL loss is calculated without the blue indicator function in Eq. 3.2 and Eq. 3.2, and (g) KCQRL w/o sol. steps, whose CL loss is calculated based only on $\mathcal{L}_{\textrm{question}_{i}}$ by ignoring $\mathcal{L}_{\textrm{step}_{i}}$. Finally, we show the question embeddings from our complete KCQRL framework.
Figure 11: Visualization of question embeddings for all ablations. For better intuition, we chose the same question from Fig. 2, and, for each of its KCs, we color the question representations sharing the same KC. Evidently, our CL loss is highly effective.
We have two main observations: (1) For the five ablations without any CL training, the representations of questions from the same KC are scattered very broadly. This highlights the importance of our CL training. (2) Our custom CL loss yields better clustered representations of questions compared to the standard CL loss by a clear margin.
Appendix H Quality of Automated KC Annotation
In this section, we provide the details how we evaluate the quality of our KCQRL’s KC annotations as part of our ablation study in Sec. 5.2.
For this, we compared our annotations against two baselines: 1) KC annotations from the original datasets and 2) KC annotations of our KCQRL without leveraging the solution steps.
@@ Appendix H Quality of Automated KC Annotation
For a structured comparison, we defined 5 criteria. Correctness is to measure which annotation has correct KC(s), relative to the other annotations. Coverage is an evaluation criterion for questions with multiple KCs. It measures how well the KCs from an annotation cover the set of KCs that the question is associated with. Specificity is a measure to compare which annotation has a more modular set of KCs rather than long and complicated ones. Of note, modular and simpler KCs enable identifying the common skills required for different questions. We additionally defined ability of integration to evaluate how well the described KCs are widely applicable to other Math problems, beyond the question being solved. Finally, overall is about choosing the best KC annotation by considering all criteria.
As explained in Sec. 5.2, we made a pair-wise comparison between KC annotations: (i) Original vs. KCQRL w/o sol. steps, (ii) Original vs. KCQRL, and (iii) KCQRL w/o sol. steps vs. KCQRL. To eliminate any potential bias from the order of KC annotations in the prompt, each time we randomly assigned KC annotations to groups A and B. Our prompt is given in Appendix H.2.
H.1 Example KC Annotations
Here, we follow our example question from Fig. 2 and provide the KC annotations of two baselines and our framework.
@@ H.1 Example KC Annotations
Original from the dataset: a) Extracting common factors of integer multiplication (ordinary type).
KCQRL w/o solution steps: a) Understanding of addition. b) Ability to perform multiplication.
Our complete KCQRL framework: a) Understanding of multiplication. b) Understanding of addition. c) Factoring out a common factor. d) Simplifications of expressions. e) Distributive property.
Takeaway: The original KC annotation is just one phrase with complex combinations of multiple KCs. It is also missing out some KCs such as addition and simplifications of expressions. On the other hand, KC annotations of KCQRL w/o solution steps are missing the important techniques asked by the problem, such as extracting the common factor. The reason is, this technique is hidden in the solution steps, which need to be provided for a better KC annotation. Overall, our complete framework provides correct set of KCs with better coverage and in a modular way in comparison to two baseline annotations.
H.2 Prompt for Quality Comparison of KC Annotations
@@
Appendix I Prompts for KC Annotation via LLMs
@@
I.1 Solution step generation
@@ I.2 KC Annotation
I.3 Solution Step-KC Mapping
As the final part, our framework maps each solution step to its associated KCs for a given question. This step is particularly needed for the Module 2 of our framework: representation learning of questions (Sec. 3.2). The prompt of this step is given below.
@@ I.3 Solution Step-KC Mapping
Appendix J Summary Statistics of KC Annotation
For XES3G5M dataset, our KCQRL framework annotated 8378 unique KCs for 7652 questions in total. Fig. 12 shows the distribution of the number KCs annotated per question. Compared to the original dataset with 1.16 KCs per question, our framework identifies 4 or 5 KCs for the majority of questions. It shows that our framework annotates the questions with more modular KCs in comparison to the original dataset.
Figure 12: Distribution of questions with different numbers of KCs annotated by our KCQRL framework for XES3G5M dataset.
Our framework identifies 2024 clusters for these 8378 unique KCs annotated. Fig. 13 shows the most frequent clusters across all the questions. To keep the plot informative, we discard the clusters that include basic arithmetic operations as they appear in the majority of the questions. To provide reproducibility and deeper insights into our KC annotations and clusters, we provide the full annotations in our repository.
Figure 13: Most frequent KCs annotated across all the questions in XES3G5M dataset. Result is shown after clustering.
Appendix K Quality of KC Annotation with Varying Sizes of LLMs
In this section, we inspect the impact of LLM’s capacity on the quality of automated KC annotation. For this, we pick different LLMs from the same family of models, namely Llama-3.2-3B, Llama-3.1-8B and Llama-3.1-70B (Dubey et al., 2024). To establish a reference annotation, we further included the KC annotations of the original XES3G5M dataset.
We followed a similar procedure to Appendix H. Specifically, we again picked 1,000 random questions from XES3G5M dataset. As we earlier found that LLMs provide better KC annotations by considering the solution steps (Sec. 5.2 and Appendix H), we provided solution steps to above Llama models for KC annotation. We used the same prompt template as in Appendix I.2. To compare four KC annotations (one original and three Llama models), we leverage GPT-4o instead of any Llama model to prevent the potential bias of Llama models towards their own generations. For comparison, we used the same five evaluation metrics defined in Appendix H), namely correctness, coverage, specificity, ability of integration and overall. We leveraged the same prompt template in Appendix H.2 by only extending it to 4 groups of KC annotations. At each inference, we randomly shuffled the order of KC annotations to avoid a potential bias from the order of KCs presented.
Table 5: Ablation study showing the relevance of automated KC annotations with varying sizes of LLMs. We report the quality (in %) for different KC annotations.
Original Llama-3.2-3B Llama-3.1-8B Llama-3.1-70B
Correctness 0.7 17.0 31.2 51.1
Coverage 1.8 19.1 47.2 31.9
Specificity 11.5 17.4 24.6 46.5
Ability of integration 3.3 17.1 37.3 42.3
Overall 1.3 17.1 36.1 45.5
Table 5 compares the quality of KC annotations of varying size of Llama models and the KC annotation of the original dataset. Specifically, it shows % of times where each model is chosen to be the best for each criterion. Overall, larger Llama models are preferred over the smaller ones. The only exception is the coverage where Llama-3.1-8B is found to be the best. After manual inspection, we found that Llama-3.1-8B generates much more KCs than the other models, which leads to increasing coverage but also decreasing correctness and specificity. Further, the overall quality of original KC annotations is only found to be best at only $\sim$1 % of the questions, when compared against three other Llama variants. This further highlights the need for designing better KC annotation mechanisms.
Appendix L Quality of KC Annotations via Human Evaluation
In this section, we conduct a human evaluation to assess the quality of KC annotations. Specifically, we randomly selected 100 questions from the XES3G5M dataset and asked seven domain expert evaluators to compare KC annotations from three sources: (i) the original dataset, (ii) Llama-3.1-70B [^8] As in Appendix K, the model was provided with solution steps to generate KC annotations, ensuring a fair comparison with our framework. (the best-performing model among Llama versions evaluated in Appendix K), and (iii) our KCQRL framework leveraging GPT-4o.
To ensure fairness, we randomly shuffled the KC annotations for each question and concealed their sources. For each question, we provided step-by-step solution steps to aid the evaluators’ assessments. While we only requested the overall preference of evaluators, we supplied detailed explanations for additional criteria (e. g., correctness, coverage, etc.) to support their decision-making. (Of note, we received the initial feedback that evaluating all five criteria for each question would be too exhaustive, so we proceeded with overall evaluations only.)
(a) % of Preferences for Different KC Annotations
(b) Majority Vote for Different KC Annotations
Figure 14: Human evaluation on KC annotations
Fig. 13(a) shows the distribution of preferences for each KC Annotation method, aggregated over all the questions and the evaluators. Only 5 % of the time the evaluators preferred the original KCs over the LLM-generated KC annotations. This percentage is even lower than our automated evaluation of KC annotations in our ablation studies (Sec. 5.2). We also found that the KC annotations of GPT-4o is preferred more than Llama-3.1-70B (55.65 % vs. 39.35 %), which confirms our choice of LLM in KC annotation module (Sec. 3.1).
Fig. 13(b) shows the number of questions for which each KC annotation method got the majority of the votes. The results are similar to the earlier distribution of preferences. The original KC annotations got the majority of the votes for only 4 questions. In comparison, for more than half of the questions, the annotations of our framework got the majority votes.
We additionally checked if there are any questions where all seven evaluators agreed on the same KC annotation. (Note that, the chance of such agreement at random is only around $\sim 0.14\,\%$.) We found that all evaluators chose GPT-4o for seven questions, and chose Llama-3.1-70B for 3 questions. On the other hand, there is no question for which all seven evaluators preferred original KC annotations. As a result, all our findings from human evaluations show that LLMs are preferred more over the KC annotations from the original dataset. Further, the strong preference over GPT-4o confirms our choice in our KC annotation module.
Appendix M Ablation Study on Different Model Embeddings
In this section, we design an ablation study to compare the quality of different question embeddings. Specifically, we get the question embeddings from Word2Vec (average over all the words in a sentence), OpenAI’s text-embedding-3-small model, and BERT (our encoder LM $E_{\psi}(\cdot)$ of representation learning module in Sec. 3.2). We follow the same procedure as we compared the question embeddings in Sec. 5.2: we get the embeddings of the question, its solution steps, and KCs concatenated as in the part (d) of Sec. 5.2. We then compare these embeddings based on the performance of downstream KT models. To establish a better reference, we further include the default performance of these KT models (i. e., with random initialized embeddings) and the performance of our KCQRL framework.
Figure 15: The quality of different embedding methods compared against our KCQRL framework.
Fig. 15 shows the performance of four KT models, AKT, DKT, DKVMN and sparseKT with five different embeddings. We have the following observations: (1) The embeddings of Word2Vec does not yield much improvement over the default version, and it might even hurt the performance. We explain this observation by that (i) Word2Vec is not advanced technique to get the semantics of a Math problem and (ii) averaging over the word embeddings can lead question embeddings to concentrate around the same region. (2) OpenAI’s text-embedding-3-small performs at a similar level to our encoder LM $E_{\psi}(\cdot)$. The reason is, although performing well on a variety of tasks, this model is not specialized in Math education. (3) Our KCQRL framework outperforms other embedding methods, which highlights the need for our representation learning module in Sec. 3.2.
Appendix N Discussion on KCQRL Usage in Real-Time
In this section, we elaborate on how our complete KCQRL framework can be deployed and maintained in real-world scenarios. Below, we discuss various situations that may arise when KCQRL is used for knowledge tracing in online learning platforms.
What happens when a new student joins the platform? This scenario has no impact on Module 1 (KC annotation via LLMs in Sec. 3.1) and Module 2 (representation learning of questions in Sec. 3.2) as the database of questions remains unchanged. KT models in our Module 3 (Sec. 3.3) are time-series models that can adapt to new sequences during inference when a new student joins. However, KT models often suffer from cold start issues (Zhang et al., 2014; Fatemi et al., 2023) when students solve only a few exercises. To address this, one can determine the minimum number of exercises required to achieve satisfactory prediction performance, as described in our multi-step ahead prediction analysis in Sec. 5.1. Students can then be guided to follow a curriculum designed by education experts until they complete enough exercises for the KT models to deliver effective predictions.
What happens when an instructor adds a new question to the platform? After training the representation learning module (Sec. 3.2), the encoder LM $E_{\psi}(\cdot)$ learns to associate questions and their solution steps with relevant KCs based on annotations provided by a more capable LLM, such as GPT-4999This can be viewed as a form of knowledge distillation from a larger LLM to a smaller one.. Consequently, when instructors add a new question to the platform, they can input the question and its solution step into the representation learning module and run the encoder LM $E_{\psi}(\cdot)$ to generate new embeddings. Details of the inference behavior of $E_{\psi}(\cdot)$ are provided in Appendix D. To evaluate student knowledge on these new questions, the downstream KT models can then run inference using the newly generated embeddings.
What happens if the LLM used for KC annotation becomes obsolete? As mentioned earlier, the LLM for KC annotation is required only during the training of the encoder LM in the representation learning module. Once the encoder LM is trained, our KCQRL framework no longer depends on the LLM from Module 1 during deployment. Therefore, even if the LLM used for KC annotation becomes obsolete, online learning platforms can continue admitting new students and allowing instructors to add new questions. These new questions can still be used to assess student knowledge effectively using our KCQRL framework.
The only exception to this scenario occurs if the platform expands its knowledge tracing to a new subject (e. g., from Math to Chemistry or Physics). In this case, KC annotation for the new question corpus will be necessary to adapt the representation learning module to the new domain via training. If the new subject is unrelated to the existing one (i. e., knowledge in one subject does not inform another), we recommend training and deploying separate branches of the KCQRL pipeline to ensure better performance for the downstream KT models.
How does KCQRL framework apply to subjects other than Math? Module 1 (KC annotation via LLMs in Sec. 3.1) in our framework can be extended to other subjects. For subjects where questions do not require multiple solution steps to arrive at the correct answer, the solution step generation part of Module 1 can be skipped, starting directly with the KC annotation process. As demonstrated in our ablation study (Sec. 5.2), our framework still outperforms the original KC annotations even without the inclusion of solution steps. Once KC annotations are obtained, the solution step–KC mapping part of Module 1 can also be omitted, as there are no solution steps to map.
In this scenario, the representation learning module (Sec. 3.2) can be trained solely using $\mathcal{L}_{\textrm{question}_{i}}$ (Eq.2). As shown in our ablation study (Sec. 5.2), this approach still improves the performance of downstream KT models. After generating the embeddings, the downstream KT model can then be trained as usual.
Finally, we emphasize that, at the time of writing, KT datasets with available question content are extremely limited; we identified only two such datasets, both in Math. We hope our work highlights the need for KT datasets in other subjects with available question corpora and inspires future research to further integrate question semantics into downstream KT models.